Continuation of this work. Yesterday, Jeff and I set up a way to have constant light exiting OFC2 to align the 1550 path for OFC3. I placed lenses according to the mode matching solution in the previous post and profiled the light coming out. Plugging in these numbers yields less than 1% mode mismatch for cavity 3. I will attempt to align 1550 light into the cavity now.